### v3

@@ 4.2 Avoidance of strictly saddle points
In this subsection, we further show that DIRL2 algorithm does not converge to fixed points that are stationary points and strict saddle points. Unlike DIRL1, the map $S$ in DIRL2 does not enjoy the property of “active manifold identification” or “model identification”. There is no guarantee that the iteration map is $C^{1}$-smooth without this property. The following additional assumption is needed to guarantee the local smoothness. One can verify that most sparse non-convex penalty in Table 1 like LPN, TAN, FRA, EXP and LOG are satisfied.
@@
$\lim\limits_{z\rightarrow 0}\frac{r^{\prime}(|z|)}{|z|}=\infty\quad\text{ and }\quad\lim\limits_{z\rightarrow 0}z\frac{r^{\prime\prime}(|z|)}{r^{\prime}(|z|)^{2}}\to 0.$
@@
Then $S_{i}^{x}(x,\epsilon)=g(z)(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x))$ with $z=\sqrt{x^{2}+\epsilon^{2}}$. We first have the following result about $g$, $S$ and $T$.
Theorem 19.
  1. (i)
    $g(z)$ is continuously differentiable on $\mathbb{R}$, so that it is Lipschitz continuous on $\mathcal{L}(F^{0})$.
  2. (ii)
    $S$ is continuously differentiable and Lipschitz continuous on $\mathcal{L}(F^{0})$ with constant $L_{S}$.
  3. (iii)
    $T$ is a lipeomorphism for $\alpha\in(0,\frac{1}{1+L_{S}})$.
Proof.
(i) It follows that
@@
by Assumption 4. Therefore, $g$ is continuously differentiable.
(ii) It is only needed to show the smoothness of $S$ when $\mathcal{J}^{*}\neq\emptyset$. We simply calculate its partial derivative and verify it is continuous at any point $(x^{*},0)\in\mathbb{R}^{2n}$.
First, consider $S_{i},i\in\mathcal{J}^{*}$. It follows from $S_{i}^{x}(x^{*},0)=0$ and $r^{\prime}(|0^{+}|)=\infty$ that
$\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}w(\delta,0)}[\delta-\tfrac{1}{\beta}\nabla_{i}f(x^{*}+\delta e_{i})]/\delta$
@@
$=$ $-\tfrac{1}{\beta}\nabla_{i}f(x^{*})\lim\limits_{\delta\rightarrow 0}g^{\prime}(0)=0.$
As for $j\neq i$, it follows that $\frac{\partial}{\partial x_{j}}S_{i}(x^{*},\epsilon^{*})=0$ and $\frac{\partial}{\partial\epsilon_{j}}S_{i}(x^{*},\epsilon^{*})=0$.
Generally, for any $i$ satisfying $z_{i}:=\sqrt{x_{i}^{2}+\delta_{i}^{2}}>0$, $S$ is smooth with respect to $x_{i}$ and $\epsilon_{i}$ with $z_{i}=\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}\neq 0$.
$\frac{\partial}{\partial x_{i}}S_{i}^{x}(x,\epsilon)=$ $\ g(z)(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x))+\tfrac{x_{i}}{z_{i}}g^{\prime}(z_{i})(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x))$ (38)
$\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x,\epsilon)=$ $\ \tfrac{\epsilon_{i}}{z_{i}}g^{\prime}(z)(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x)),$ (39)
$\frac{\partial}{\partial x_{j}}S_{i}^{x}(x,\epsilon)=$ $\ -g(z)\tfrac{1}{\beta}\nabla^{2}_{ij}f(x),\quad j\neq i$ (40)
$\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x,\epsilon)=$ $\ 0,\quad j\neq i.$ (41)
Now for any $i\in\mathcal{J}^{*}$, as $z_{i}\to 0$, it follows from $\frac{x_{i}}{z_{i}}\leq 1$ and $zr^{\prime\prime}(z)/(r^{\prime}(z))^{2}\to 0$ that
$\lim_{z_{i}\to 0}\frac{\partial}{\partial x_{i}}S_{i}^{x}(x,\epsilon)=0=\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*}),\ \lim_{z_{i}\to 0}\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x,\epsilon)=0=\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x^{*},\epsilon^{*})$
$\lim_{z_{i}\to 0}\frac{\partial}{\partial x_{j}}S_{i}^{x}(x,\epsilon)=0=\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*}),\ \lim_{z_{i}\to 0}\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x,\epsilon)=0=\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x^{*},\epsilon^{*})$
Overall, we have shown that $S$ is continuously differentiable. The Lipschitz continuity of $S$ then trivially comes from (i) and the boundedness of $\mathcal{L}(F^{0})$.
(iii) This is straightforward by applying [23, Lemma 2.13 and 2.14]. ∎
@@
(Unstable Fixed Point) Suppose $x^{*}$ is a stationary point of (P). If $x^{*}$ is a strict saddle point of (P), then the corresponding $(x^{*},\epsilon^{*})$ with $\epsilon^{*}=0$ is an unstable fixed point of $T$. Moreover, for $\alpha<\beta/\rho$, $DT(x^{*},0)$ is invertible.
@@
Consider the Jacobian of $S$ at a fixed point $(x^{*},0)$. Define $H\in\mathbb{R}^{|\mathcal{I}^{*}|\times n}$ as follows:
$\forall i\in\mathcal{I}^{*},\ h_{ii}=$ $\ \frac{(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}-\frac{(|x_{i}^{*}|+\tfrac{\lambda}{\beta}r^{\prime}(x^{*}_{i}))\tfrac{\lambda}{\beta}\left(|x^{*}_{i}|r^{\prime\prime}(|x^{*}_{i}|)-r^{\prime}(|x^{*}_{i}|)\right)}{\left(|x^{*}_{i}|+\tfrac{\lambda}{\beta}r^{\prime}(|x^{*}_{i}|)\right)^{2}},$
$=$ $\ \frac{(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}-\frac{\tfrac{\lambda}{\beta}\left(r^{\prime\prime}(|x^{*}_{i}|)-\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}\right)}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}$
$=$ $\ \frac{1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*})-\tfrac{\lambda}{\beta}r^{\prime\prime}(|x^{*}_{i}|)+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}$
$=$ $\ 1-\tfrac{1}{\beta}\frac{\nabla^{2}_{ii}f(x^{*})+\lambda r^{\prime\prime}(|x^{*}_{i}|)}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}$
$\forall j\neq i,\ h_{ij}:=$ $\ \frac{\partial}{\partial x_{j}}S_{i}(x^{*},0)=-\tfrac{1}{\beta}\frac{\nabla^{2}_{ij}f(x^{*})}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|x_{i}^{*}|)}{|x_{i}^{*}|}},$
by (38) and (40). Therefore, $H_{\mathcal{I}^{*}\mathcal{I}^{*}}=I-\tfrac{1}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$,
$DS(x^{*},0)=$ $\begin{bmatrix}H_{\mathcal{I}^{*}\mathcal{I}^{*}}&H_{\mathcal{I}^{*}\mathcal{J}^{*}}&0\\ 0&0&0\\ 0&0&\mu I\end{bmatrix}$
$DT(x^{*},0)=$ $\begin{bmatrix}(1-\alpha)I+\alpha H_{\mathcal{I}^{*}\mathcal{I}^{*}}&H_{\mathcal{I}^{*}\mathcal{J}^{*}}&0\\ 0&(1-\alpha)I&0\\ 0&0&(1-\alpha(1-\mu))I\end{bmatrix}.$
Hence, the eigenvalues of $DS(x^{*},0)$ are the same as
@@
where $W^{*}:=W(x^{*},0)$ and $W^{*}_{ii}<\infty,i\in\mathcal{I}^{*}$. If $\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0$, it follows from Lemma [49, Proposition 6.1] that $\lambda_{\min}((I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0$. Hence, $DT(x^{*},0)$ has an eigenvalue greater than 1, meaning $(x^{*},0)$ is the unstable fixed point of $T$. Moreover, since $\|\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\|\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\|\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\rho$, it is clear that for $\alpha<\beta/\rho$, $DT(x^{*},0)$ is invertible. ∎
Similarly, we can derive the global escape property of ${\rm DIRL}_{2}$ algorithm.
@@
It follows directly from Theorem 19 and 20 that 1) map $T$ is local $C^{1}$-smooth around any stationary points; 2) all stationary points $x^{*}$ are unstable fixed points; 3) $T$ is a lipeomorphism. Then by Corollary 1, the set of initial conditions attracted by such fixed points has zero Lebesgue measure. ∎

### v4

@@ 4.2 Avoidance of strictly saddle points
In this subsection, we further show that the DIRL2 algorithm does not converge to fixed points that are stationary points and strict saddle points. Unlike DIRL1, the map $S$ in DIRL2 does not enjoy the property of “active manifold identification” or “model identification”. There is no guarantee that the iteration map is $C^{1}$-smooth without this property. The following additional assumption is needed to guarantee the local smoothness. One can verify that most sparse nonconvex penalty in Table 1 like LPN, TAN, FRA, EXP and LOG are satisfied.
@@
$\lim\limits_{z\rightarrow 0}r^{\prime}(|z|)=\infty\ \text{ and }\ \lim\limits_{z\rightarrow 0}z\frac{r^{\prime\prime}(|z|)}{r^{\prime}(|z|)^{2}}\to 0.$
@@
We can write $S_{i}^{x}(x,\epsilon)=g(z)(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x))$ with $z=\sqrt{x^{2}+\epsilon^{2}}$. It follows that
@@
by Assumption 4. Therefore, $g$ is continuously differentiable in $\mathbb{R}$.
We have the following result for $S$ and $T$.
Theorem 19.
  1. (i)
    $S$ is $C^{1}$ smooth and Lipschitz continuous on $\mathcal{L}(F^{0})$ with constant $L_{S}$; moreover
    $DS(x^{*},\bm{0})=\ \begin{bmatrix}H_{\mathcal{I}^{*}\mathcal{I}^{*}}&H_{\mathcal{I}^{*}\mathcal{J}^{*}}&,\bm{0}\\ 0&0&0\\ 0&0&\mu I\end{bmatrix},$ (40)
    where $H_{\mathcal{I}^{*}\mathcal{I}^{*}}=I-\tfrac{1}{\beta}(I+\tfrac{\lambda}{\beta}W_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$.
  2. (ii)
    $T$ is a lipeomorphism for $\alpha\in(0,\frac{1}{1+L_{S}})$ with
    $DT(x^{*},\bm{0})=\ \begin{bmatrix}(1-\alpha)I+\alpha H_{\mathcal{I}^{*}\mathcal{I}^{*}}&H_{\mathcal{I}^{*}\mathcal{J}^{*}}&\bm{0}\\ \bm{0}&(1-\alpha)I&\bm{0}\\ \bm{0}&\bm{0}&(1-\alpha(1-\mu))I\end{bmatrix}.$
Proof.
(i) We simply calculate its partial derivative and verify that it is continuous at any point $(x^{*},\epsilon^{*})\in\mathbb{R}^{2n}$.
Consider $i$ satisfying $z_{i}^{*}:=\sqrt{(x_{i}^{*})^{2}+(\epsilon_{i}^{*})^{2}}>0$. In this case, $S$ is smooth with respect to $x_{i}$ and $\epsilon_{i}$, and
$\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ g(z^{*}_{i})(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))+\tfrac{x_{i}^{*}}{z_{i}^{*}}g^{\prime}(z_{i}^{*})(x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*}));$ (41)
$\frac{\partial}{\partial x_{j}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ -g(z^{*}_{i})\tfrac{1}{\beta}\nabla^{2}_{ij}f(x^{*}),\quad j\neq i;$ (42)
$\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x^{*},\epsilon^{*})=$ $\ \tfrac{\epsilon_{i}^{*}}{z_{i}^{*}}g^{\prime}(z^{*}_{i})(x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*}));$ (43)
$\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x^{*},\epsilon^{*})=$ $\ 0,\quad j\neq i.$ (44)
Consider $i$ satisfying $x_{i}^{*}=0$ and $\epsilon_{i}^{*}=0$. It follows from $S_{i}^{x}(x^{*},\bm{0})=0$ and $r^{\prime}(|0^{+}|)=\infty$ that
$\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}w(\delta,0)}[\delta-\tfrac{1}{\beta}\nabla_{i}f(x^{*}+\delta e_{i})]/\delta$ (45)
@@
$=$ $-\tfrac{1}{\beta}\nabla_{i}f(x^{*})\lim\limits_{\delta\rightarrow 0}g^{\prime}(0)=0,$
$\frac{\partial}{\partial x_{j}}S_{i}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}w(\delta,0)}[\delta-\tfrac{1}{\beta}\nabla_{i}f(x^{*}+\delta e_{j})]/\delta$ (46)
$=$ $\ -\tfrac{1}{\beta}\nabla_{i}f(x^{*})\lim\limits_{\delta\rightarrow 0}[g(\delta)-g(0)]/\delta=0,\quad j\neq i;$
$\frac{\partial}{\partial\epsilon_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}w(0,\delta)}[-\tfrac{1}{\beta}\nabla_{i}f(x^{*})]/\delta$ (47)
$=$ $\ -\tfrac{1}{\beta}\nabla_{i}f(x^{*})\lim\limits_{\delta\rightarrow 0}[g(\delta)-g(0)]/\delta=0;$
[…]
Now we verify that the Jacobian is continuous at $(x^{*},\epsilon^{*})$. It suffices to show this for any $i$ satisfying $x_{i}^{*}=0$ and $\epsilon_{i}^{*}=0$; indeed, it holds that It follows from (41), the fact that $\frac{x_{i}}{z_{i}}\leq 1$ and $g(0)=0,g^{\prime}(0)=0$ that
$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial x_{i}}S_{i}^{x}(x,\epsilon)=0=\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*}).$
It follows from (42) and $g(0)=0$ that
$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial x_{j}}S_{i}^{x}(x,\epsilon)=0=\frac{\partial}{\partial x_{j}}S_{i}^{x}(x^{*},\epsilon^{*}).$
It follows from (43), the fact that $\frac{x_{i}}{z_{i}}\leq 1$ and $g^{\prime}(0)=0$ that
$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x,\epsilon)=0=\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x^{*},\epsilon^{*}).$
It follows from (44) that
$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x,\epsilon)=0=\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x^{*},\epsilon^{*}).$
Overall, we have shown that $S$ is continuously differentiable.
To derive the matrix form of the Jacobian, it follows from (41) that for $i\in\mathcal{I}^{*}$,
$h_{ii}:=$ $\ \frac{\partial}{\partial x_{j}}S_{i}(x^{*},\bm{0})$
$=$ $\ \frac{(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}-\frac{(|x^{*}_{i}|+\tfrac{\lambda}{\beta}r^{\prime}(x^{*}_{i}))\tfrac{\lambda}{\beta}\left(|x^{*}_{i}|r^{\prime\prime}(|x^{*}_{i}|)-r^{\prime}(|x^{*}_{i}|)\right)}{\left(|x^{*}_{i}|+\tfrac{\lambda}{\beta}r^{\prime}(|x^{*}_{i}|)\right)^{2}},$
$=$ $\ \frac{(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}-\frac{\tfrac{\lambda}{\beta}\left(r^{\prime\prime}(|x^{*}_{i}|)-\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}\right)}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x_{i}^{*}|}}$
$=$ $\ \frac{1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*})-\tfrac{\lambda}{\beta}r^{\prime\prime}(|x_{i}^{*}|)+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}$
$=$ $\ 1-\tfrac{1}{\beta}\frac{\nabla^{2}_{ii}f(x^{*})+\lambda r^{\prime\prime}(|x_{i}^{*}|)}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}$
It follows from (42) that for $j\neq i$,
$h_{ij}:=\frac{\partial}{\partial x_{j}}S_{i}(x^{*},\bm{0})=-\tfrac{1}{\beta}\frac{\nabla^{2}_{ij}f(x^{*})}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}.$
We then combine this with (43)–(48) to derive (40) of the Jacobian $DS(x^{*},\bm{0})$. The Lipschitz continuity of $S$ trivially leads to the boundedness of $\mathcal{L}(F^{0})$.
(ii) This is straightforward by applying [23, Lemma 2.13 and 2.14]. The Jacobian $DT(x^{*},\bm{0})$ is obtained directly by (i). ∎
@@
(Unstable Fixed Point) Suppose Assumption 1-4 hold. Let $x^{*}$ be a stationary point of (P). If $x^{*}$ is a strict saddle point of (P), then the corresponding $(x^{*},\epsilon^{*})$ with $\epsilon^{*}=\bm{0}$ is an unstable fixed point of $T$. Moreover, for $\alpha<\beta/\rho$, $DT(x^{*},\bm{0})$ is invertible.
@@
It follows from Theorem 19(ii) that we only have to check the eigenvalues of
@@
where we have $W^{*}_{ii}=w(x_{i}^{*},0)<\infty,i\in\mathcal{I}^{*}$. If $\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0$, it follows from Lemma [49, Proposition 6.1] that $\lambda_{\min}((I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0$. Hence, $DT(x^{*},\bm{0})$ has an eigenvalue greater than 1, meaning $(x^{*},\bm{0})$ is the unstable fixed point of $T$. Moreover, since $\|\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\|\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\|\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\rho$, it is clear that for $\alpha<\beta/\rho$, $DT(x^{*},\bm{0})$ is invertible. ∎
Now we can derive the global escape property of ${\rm DIRL}_{2}$ algorithm.
@@
It follows directly from Theorem 19 and 20 that: (i) map $T$ is local $C^{1}$-smooth around any stationary points; (ii) all stationary points $x^{*}$ are unstable fixed points; (iii) $T$ is a lipeomorphism. Then by Corollary 1, the set of initial conditions attracted by such fixed points has zero Lebesgue measure. ∎